Sunil,who holds a bill worth $40000,gives the bill to a bank with instructions to accept payment at maturity.If the bank charges 1% commission,Find the amount received by sunil
1
Expert's answer
2020-06-02T19:13:46-0400

Sunil will recieve $ 40000 except for 1% comission, hence

"40000-40000*0.01=40000*0.99=39600" dollars